OpenStudy (anonymous):

help please Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain. a_n = 40 + 7(n − 1) where n ≥ 40 a_n = 40 + 7(n − 1) where n ≥ 1 a_n = 40 − 7(n − 1) where n ≥ 40 a_n = 40 − 7(n − 1) where n ≥ 1

OpenStudy (anonymous):

n a_n 1 40 2 47 3 54

OpenStudy (anonymous):

a_n = 40 + 7(n − 1) where n ≥ 1

OpenStudy (anonymous):

why xo?

OpenStudy (anonymous):

how did you get that and because it was my email so they made that myuser

OpenStudy (anonymous):

the table is n a_n 1 40 2 47 3 54 the domain is the values for "n" notice that they start from 1 upwards, hence we say n>=1 correct?

OpenStudy (anonymous):

you are given 4 answers, which of them matches the a_n values from the table above?

OpenStudy (anonymous):

Okay but how do you know the +7 part sorry I'm not good at this part of math

OpenStudy (anonymous):

oh +7 is how much it goes up

OpenStudy (anonymous):

when n = 1 the series is 40 and a_1 = 40 + 7(1 − 1) = 40 and so on for the remaining terms 2, 3, 4, ... xo